Astroemeria Berhad has a capital structure consisting of 30% debt and 70% equity. Astroemeria Berhad debt currently has an 8% yield to maturity. The risk-free rate (krf) is 5%, and the market risk premium (km-krf) is 6%. Using the CAPM, Alstroemeria Berhad estimates that the cost of equity is currently 12%. The company has a 28% tax rate.

Required:
i. Calculate Alstroemeria Berhad current Weighted Average Cost of Capital (WACC). (3 Marks)

ii. What is the current beta (ß) on Astroemeria Berhad common stock? (2 Marks)

iii. Alstroemeria Berhad is considering changing its capital structure to 40% debt and 60% equity. If the company went ahead with the proposed change, the yield to maturity of the company's bond would rise to 9.5%; and the beta would also change to 1.6. Should the company proceed with the restructure? What would you advise Astroemeria Berhad? (4 Marks)
Answer #1

Step 1

For calculating average cost of capital, first we have to calculate cost of capital on debt as well as equity.

In the question it has been given that debt and equity are 30% and 70%. We may assume that total capital is $1 million, Then Debt is $ 300,000 and equity is $ 700,000.

Step 2

We have been given that debt currently has 8% yield and cost of equity is 12%.

The WACC would be = E/V X Re + D/V X Rd X (1-Tc)

Where E is market value of equity

V is total capital

D is Market value of Debt

Re is cost of equity

Rd is cost of Debt and

Tc is Tax Rate.

Step 3

Thus WACC would be 700,000/1000,000 X 0.12 + (300,000/1000,000 X 0.08) X (1 - 0.28)

= 0.084 + 0.024 X 0.72 = 0.084 + 0.01728 = 0.04128

Thus WACC is 4.128 %.

Current beta

Step 4

Beta = (Ra - Krf) / (Km -Krf)

Where Ra is return on capital

Krf is Risk free rate

Km is expected rate of market

Step 5

We have been given that Risk free rate is 5%, market risk premium is 6%. Expected return on security is 700,000/1000,000 X 0.12 + 300,000/1000,000 X 0.08 = 0.084 + 0.024 = 0.108

Step 6

Thus Beta =

(0.108 -0.05 )/ (0.06 - 0.05) = (0.058)/(0.01) =5.8.